# Explainable Machine Learning Identification of Superconductivity from Single-Particle Spectral Functions

### v1

@@ Abstract
Beyond classifying superconductivity from ARPES spectra, we also aim to extract interpretable physical parameters from the ML model to deepen our understanding of quantum materials. Specifically, we seek to identify which single-particle spectral features are linked to the superconductivity long-range order via an occlusion-based analysis for the ML model [22]. This analysis measures the significance of specific portions of an ARPES spectrum to the model’s classification decision by assessing the saliency, defined as Eq. (9) in Methods. We analyze saliency for each individual energy level, which reflects the change in the phase prediction, in terms of $\bar{p}_{\textrm{SC}}$ when an entire row of the spectrum is obscured or occluded.
@@
As shown in Fig. 5, we obtain the saliency as a function of binding energy for each spectrum in both BSCCO OD50 and OD58 experimental datasets. A notable feature across various spectral samples is a pronounced peak at $\omega=0$. Remarkably, the saliency magnitude significantly decreases for spectra acquired at temperatures above $T_{c}$. This observation suggests that the ML model identifies the nuanced distribution of spectral weight near and above the gap center as the key for pinpointing the true $T_{c}$. The spectral weight just above the gap center is particularly sensitive to the upper branch formation of the Bogoliubov quasiparticles, whose clear separation from the lower branch is a major empirical identifier of superconducting phase ordering in cuprates [8]. In an intriguing parallel, a recent study inferred electronic entropy from a continuous sequence of ARPES spectra, achieved through meticulous temperature and energy calibration [15]. This study deduced that the temperature-derivative of weighted spectral intensity near $\omega=0$ could effectively determine the $T_{c}$. Our ML model’s explanation aligns with this conclusion. However, different from the temperature-derivative approach, the ML classification and the saliency analysis operate on each individual ARPES snapshot, without relying on temperature and phase information. The experimental data, used solely for domain alignment, remain unlabeled. This approach of unsorted-snapshot-based classification approach is particularly valuable for general material design, where precise parameter tuning and sequential measurements are impractical, such as under extreme or non-equilibrium conditions.
Recalling the original motivation for this spectrum-based phase classification, we must particularly investigate the fluctuating superconducting (fluc-SC) regime in the normal state. This regime is characterized by prominent short-range Cooper pair fluctuations and the single-particle gap comparable to that observed below $T_{c}$[5, 6, 7, 8]. To further assess the robustness of our model, we expand the classification to include three states: SC, fluc-SC, and gapless non-SC states. As detailed in Supplementary Note 3, the spectra stemming from SC and gapless non-SC states are still correctly classified under this ternary classification. However, distinguishing the fluc-SC phase proves challenging, resulting in an overall accuracy of 85.4% and 75.6% for BSCCO OD50 and OD58 samples, respectively. Despite the reduced accuracy, domain adaptation remains crucial in the ternary classification. The domain classifier in DANN improves the accuracy compared to traditional CNN by more than 10% and 30% with (Supplementary Table 1) and without (Supplementary Table 2) the ensemble average, respectively. An interesting observation is that the majority (15 out of 17) of (gapped) fluc-SC states are misclassified as gapless non-SC states. This misclassification occurs regardless of the gap, which was typically read by APRES at first glance and regarded as a fingerprint for long-range orders. Instead, our ternary classification results indicate that traditional single-particle spectroscopy interpretations may oversimplify the complexities inherent in quantum materials. Distinguishing the gapped fluc-SC state from the gapless non-SC state is more challenging than distinguishing between the equally gapped SC and fluc-SC states. This outcome also indicates that the ML model correctly recognizes the intrinsic similarities between fluc-SC and non-SC states, as both belong to the normal phase of these materials.
@@
In summary, this study demonstrates that a transferable ML model trained without labeled experimental data can successfully classify the superconductivity of quantum materials based on real experimental ARPES spectra. This achievement paves the way for using single-particle spectra to identify long-range orders, even when the deterministic features do not conform to traditional gap analysis. The transferability of our model is demonstrated by its application to ARPES spectral datasets collected from two materials with distinct compositions measured at various temperatures. The snapshot-based classification without the need to track temperature-dependent trends enables high-throughput characterization of superconductivity in correlated materials, with potential in pump-probe spectra where phases lose coherence [23, 24, 25]. Furthermore, the physical intuition extracted from the ML model using the occlusion-based explanation method aligns well with domain knowledge. Our approach fits within the growing emphasis on transfer learning [26, 27] and domain adaptation techniques [28, 29] in scientific applications to overcome data scarcity. Despite the challenges compared to existing efforts (fewer available experimental data, a larger discrepancy between simulation and experiment, and more complicated underlying patterns), our model achieves satisfactory performance in classifying superconductivity in quantum materials, thereby enabling high-throughput material design and synthesis experiments that require immediate characterizations [30].
Methods
.1 Simulated ARPES Data
ARPES data simulation is primarily based on the phenomenological model described in Ref. 8. The superconducting pairing onset temperature $T_{\rm gap}$ is determined by the gap opening temperature, and the thermodynamic transition temperature $T_{c}$ is determined by the steepest change of quasiparticle linewidth [20, 15]. The spectral function used in the simulation takes the following form:
@@ Methods
where $E_{\mathbf{k}}=\sqrt{\epsilon_{\mathbf{k}}^{2}+\Delta_{\mathbf{k}}^{2}}$ is the Bogoliubov quasiparticle dispersion. The superconducting gap $\Delta_{\mathbf{k}}$ has the following approximated temperature dependence:
@@
States between $T_{c}$ and $T_{\text{gap}}$ are labeled as fluc-SC. The quasiparticle width $\sigma(T)$ is modelled by a rounded step function plus a Fermi liquid self energy that mimics realistic experimental values in units of eV and Kelvin:
@@
Energy-momentum resolution is treated as a 2D Gaussian convolution on the 2D spectral function, where the momentum resolution is taken as the typical 10-pixel size. The background is treated as 10 times the average per-pixel intensity across the simulated cut multiplied by the Fermi function, which mimics the momentum-scrambling secondary scattering process in photoemission. Since we are mostly in intermediate to high count rate regime, the noise distribution is approximated as a Gaussian with a standard deviation of $\alpha\sqrt{N_{ij}}$, where $N_{ij}$ is the simulated count at pixel $\{i,j\}$, and $\alpha$ is an input parameter to control the signal to noise ratio. This approximation mimics the Poisson noise observed in real experiments, which is generally in the high-count-rate regime and asymptotically approaches a Gaussian noise. The entire spectrum takes an absolute value to eliminate negative counts due to the application of Gaussian noise at extremely low count regimes, which has negligible impact on the main spectral region of interest. A comparison of the parameter space of simulated and experimental data can be found in Supplementary Note 4. A total of 1,745 simulated spectra were obtained, with 80% (1,395) being utilized as training data.
.2 Domain adaptation
@@
DANN is a domain adaptation technique designed to reduce the domain distribution shift between the source (simulated) and target (experimental) domains. The model architecture comprises three parts as shown in Fig. 2: the feature extractor ($G_{f}(\cdot;\theta_{f})$) with parameters $\theta_{f}$, the label predictor ($G_{y}(\cdot;\theta_{y}$)) with parameters $\theta_{y}$, and the domain classifier ($G_{d}(\cdot;\theta_{d}$)) with parameters $\theta_{d}$. Training DANN involves minimizing the classification loss of the class classifier using labels from the source domain while simultaneously maximizing the error of the domain classifier using a gradient reversal layer (GRL) (denoted as $\mathcal{R}(\cdot)$). We define the $\mathcal{L}_{y}$ and $\mathcal{L}_{d}$ as the corresponding cross-entropy loss [31] of label prediction and domain classification. The GRL acts as an identity transformation during forward propagation and flips the sign of $\partial\mathcal{L}_{d}/\partial\theta_{f}$ during backpropagation. The training objective is formulated as follows:
$E(\theta_{f},\theta_{y},\theta_{d})=\mathbb{E}_{(\mathbf{x}_{i}^{s},y_{i}^{s})\sim\mathcal{D}_{s}}\mathcal{L}_{y}\big{(}G_{y}(G_{f}(\textbf{x}_{i}^{s};\theta_{f});\theta_{y}),y_{i}^{s}\big{)}\\ -\lambda\Big{(}\mathbb{E}_{\mathbf{x}_{i}^{s}\sim\mathcal{D}_{s}}\mathcal{L}_{d}\big{(}G_{d}(\mathcal{R}(G_{f}(\textbf{x}_{i}^{s};\theta_{f}));\theta_{d}),d^{s}\big{)}\\ +\mathbb{E}_{\mathbf{x}_{j}^{t}\sim\mathcal{D}_{t}}\mathcal{L}_{d}\big{(}G_{d}(\mathcal{R}(G_{f}(\textbf{x}_{j}^{t};\theta_{f}));\theta_{d}),d^{t}\big{)}\Big{)}\,,\end{aligned}$ (4)
@@
where $\mathbb{E}$ represents the expected value of the set; $\mathbf{x}_{i}^{s}$ denotes a simulated spectrum with phase label $y_{i}^{s}$ from the source domain $D_{s}=\{(\mathbf{x}_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}}$; $\mathbf{x}_{j}^{t}$ denotes an experimental spectrum in the target domain $D_{t}=\{(\mathbf{x}_{j}^{t})\}_{j=1}^{n_{t}}$; $d^{s}$ and $d^{t}$ are the domain labels. Here, $\lambda$ is the adaptation parameter, balancing the two objectives that shape the features during training. Upon training completion, the label predictor is capable of predicting labels for both source and target domain samples.
.3 Model architecture and implementation
@@
Our optimized model employs four convolutional layers with hidden channels of 16, 32, 64, and 64, respectively. Each layer utilizes a 3$\times$3 convolutional kernel, a stride of 1, and padding of 2, followed by max pooling with the same kernel size and a stride of 2, and an activation function. All the activation functions used in this network are rectified linear units (ReLU) except for the first convolutional layer. To alleviate the dying ReLU problem where some neurons become permanently inactive and only output 0 for any input during training, the first convolutional layer ReLU was replaced with the leaky ReLU (LeakyReLU) [32] using a default constant slope of 0.01. Post-convolution, the network applies adaptive average pooling to the feature maps, which are then flattened to interface with the subsequent fully connected layers. The domain classifier and label predictor are both two-layer fully connected neural networks with 64 hidden dimensions each. All models are implemented using PyTorch 1.12.1. The optimized training procedure employs the Adam optimizer [33] with a learning rate of 0.0005, weight decay (L2 penalty, 0.001), and early stopping. Dropout with a probability of 0.4 is applied after the last convolutional layer and in every layer of each fully connected neural network. Training is performed with mini-batches of 4 spectra for 150 epochs. The adaptation parameter is set to be 1.2 and is gradually increased from 0 through:
@@
.4 Model evaluation and hyperparameter tuning
Evaluating the performance of an unsupervised domain adaptation (UDA) model is challenging due to the lack of the target domain labels [34, 35, 36]. As the domain classifier acts as a regularizer on CNN [19], we first optimize the CNN model without the domain classifier using simulated ARPES data (detailed in Supplementary Note 5). We then fine-tune the adaptation parameter based on the transfer score (TS) metrics [36], which evaluates the transferability and discriminability of the feature space. The TS is formulated as follows:

### v2

@@ II.3 Physical explanation of the ML model
Beyond classifying SC from ARPES spectra, we also aim to extract interpretable physical intuitions from the ML model to deepen our understanding of quantum materials and, especially, to identify single-particle spectral features directly linked to a thermodynamic phase transition. Specifically, we exploit an occlusion-based analysis for the ML model to identify which single-particle spectral features are linked to the SC long-range order [22]. The occlusion-based analysis measures the impact of a specific data feature on a model’s output when that feature gets blocked. In the context of analyzing ARPES spectra, the occlusion is realized by blocking a region of the spectral function information via an energy-resolved occluding patch.
The occluded spectrum, $\tilde{A}(\mathbf{k},\omega;\nu)$, of the original spectrum, $A(\mathbf{k},\omega)$, is defined as
Here, $A_{0}(\mathbf{k},\nu)$ is a single baseline spectrum at energy $\nu$, which provides the spectral function value in the occluded region. Following recent feature attribution studies [23, 24], we set $A_{0}(\mathbf{k},\nu)$ as the training distribution to ensure robust explanation performance across different pixel intensities (see details in supplemental methods and Figure S4).
The change in the predicted probability before and after occlusion is usually referred to as saliency. By shifting the occluding patch in Eq. (1) along the energy axis, we can determine the saliency for each spectrum sample as a function of energy $\nu$:
where $\bar{p}_{\textrm{SC}}[\cdot]$ represents the ML model mapping from an input ARPES spectrum to the predicted SC probability $\bar{p}_{\textrm{SC}}$. To focus on features that positively contribute to the classification, negative saliency values are filtered out by a rectified linear unit (ReLU) function [25] (the max operation in Eq.2).
@@
As shown in Fig. 5, we obtain the saliency as a function of binding energy for each spectrum in both BSCCO OD50 and OD58 experimental datasets. A notable feature across various spectral samples is a pronounced peak at zero energy ($\nu=0$). Remarkably, the saliency magnitude significantly decreases for spectra acquired at temperatures above $T_{c}$. This observation suggests that the ML model identifies the nuanced distribution of spectral weight near and above the gap center as the key for pinpointing the true $T_{c}$. The spectral weight just above the gap center is particularly sensitive to the upper branch formation of the Bogoliubov quasiparticles, whose clear separation from the lower branch is a major empirical identifier of superconducting phase ordering in cuprates [5]. In an intriguing parallel, a recent study inferred electronic entropy from a continuous sequence of ARPES spectra, achieved through meticulous temperature and energy calibration [6]. This study deduced that the temperature-derivative of weighted in-gap spectral intensity near the Fermi level could effectively determine the $T_{c}$. Our ML model’s explanation aligns with this conclusion, and also naturally points to the minimal experimental requirement for our method to work: enough resolution and signal-to-noise to resolve in-gap spectral weight. However, different from the temperature-derivative approach, the ML classification and the saliency analysis operate on each individual spectrum, without relying on temperature dependent information. The experimental data, used solely for domain alignment, remain unlabeled. This approach of single-snapshot-based classification approach is particularly valuable for general material design, where exhaustive parameter tuning and sequential measurements are impractical, such as under extreme or non-equilibrium conditions.
II.4 Ternary state classification
We now turn to investigate the fluctuating superconducting (fluc-SC) regime within the normal state. This regime is characterized by pronounced short-range Cooper pair fluctuations, manifesting as a single-particle gap comparable to that observed below $T_{c}$[4, 12, 13, 5]. To further assess the robustness of our model and method, we expand the classification task to include three states: SC, fluc-SC, and gapless non-SC states. Specifically, we re-label simulated spectral data to reflect these three states and retrain the DANN model, while keeping the experimental data unlabeled. This model is then applied to individual ARPES experimental spectra to distinguish whether the corresponding sample material is in the SC, fluc-SC, or gapless non-SC state.
Figure 6: Ternary classification results for two BSCCO samples. A Predicted probability in the ternary graph for superconducting (red), fluctuating-superconducting (green), and gapless non-superconducting (blue) phases using the DANN model for experimental spectral data measured in the BSCCO OD50 sample. The proximity of each point to each vertex represents the probability for the corresponding phase, while the color reflects its true label determined by experiments. B Confusion matrix for the ternary classification in A. C,D Same as A and B but for BSCCO OD58 samples.
As shown in Fig. 6, the spectra stemming from SC and gapless non-SC states are still correctly classified under this ternary classification. However, distinguishing the fluc-SC phase presents greater challenges, resulting in an overall accuracy of 85.4% for the BSCCO OD50 sample and 75.6% for the OD58 sample. An interesting observation is that the majority (15 out of 17) of gapped fluc-SC states are misclassified as gapless non-SC, despite the presence of a gap, which is traditionally considered a fingerprint of long-range order. Instead, our ternary classification results correctly identify the inherent similarities between gapped fluc-SC and gapless non-SC states, both of which exist above $T_{c}$ in the normal state. This result suggests that traditional single-particle spectroscopy, which relies heavily on gap detection, oversimplifies the complexities fluctuating states to a single gap value. The gapped fluc-SC state and the gapless non-SC state are inherently a crossover, which also proves more difficult to distinguish compared to the detection of true SC long-range (phase) ordering. Although the accuracy decreases for the fluc-SC phase, domain adaptation remains crucial in improving performance. It enhances the accuracy of the phase classification by more than 10% and 30% compared to a traditional CNN, with (Table S1) and without (Table S2) ensemble averaging, respectively.
III Discussion
@@
Our study demonstrates that a transferable ML model, trained without labeled experimental data, effectively classifies the SC of strongly correlated materials based on single-snapshot ARPES spectrum. This breakthrough suggests a powerful new tool for identifying thermodynamic phase transitions and long-range orders from electronic spectroscopy, even when the deterministic features do not conform to traditional gap analysis. The model’s transferability between simulation and experiment is demonstrated through its accurate classification of ARPES spectra from two materials with distinct compositions, gap sizes, and bare band dispersions and that are measured at a variety of temperatures. This pre-trained model is also expected to accurately classify other superconductive materials similar to cuprates in a traditional inductive manner (see Note S3 and Tables S3 and S4 for details). For materials significantly different from cuprates, accurate classifications can still be achieved through transductive learning by fine-tuning the model with corresponding unlabeled experimental spectra (Note S3). In both cases, our model bypasses the need for tracking temperature-dependent trends, thereby enabling non-contact, in situ, and operando characterization of superconductivity in correlated materials, with potential in pump-probe spectra where phases lose coherence [26, 27, 28]. Furthermore, the physical intuition gained from the ML model, revealed through occlusion-based attribution analysis, closely matches the recent conclusion from temperature-dependent analysis [6]. Our approach fits into the growing emphasis on transfer learning [29, 30] and domain adaptation techniques [31, 32] in scientific applications to overcome data scarcity. Despite the unique challenges compared to existing efforts, including fewer available experimental data, a larger discrepancy between simulation and experiment, and more complicated underlying patterns, our model achieves impressive results in classifying superconductivity in quantum materials, with the potential to classify other thermodynamic phase transitions after retraining with new simulated data generated from relevant phenomenological equations. Our model paves the way for high-throughput material design and synthesis experiments that require immediate characterizations [33].
IV Methods
IV.1 Simulated ARPES data
ARPES data simulations are based on the phenomenological model described in He et al. [5], which applies to a broad range of superconducting materials [34, 16, 35, 36]. The spectral function used in the simulation is given by:
@@ IV.1 Simulated ARPES data
where $E_{\mathbf{k}}=\sqrt{\epsilon_{\mathbf{k}}^{2}+\Delta_{\mathbf{k}}^{2}}$ is the Bogoliubov quasiparticle dispersion, $\Delta_{\mathbf{k}}$ is the momentum dependent superconductivity order parameter, and $\sigma$ is the spectral broadening.
The superconducting order parameter $\Delta_{\mathbf{k}}$ is determined by the general band BCS gap equation, with an approximate closed-form solution over the entire temperature range:
@@
Here, $T_{\rm gap}$ denotes the superconducting pairing onset temperature, while $T_{c}$ is the thermodynamic transition temperature. Both parameters are randomly sampled in the simulations to enhance model transferability (see Note S4 and Figure S2) [20, 6]. States between $T_{c}$ and $T_{\text{gap}}$ are labeled as fluc-SC. The coefficient 2.34 in Eq. (4) is specific to $d$-wave superconductors and may slightly vary for different Fermi-surface geometries [37]. However, as it simply rescales temperature, which remains hidden in DANN training, its value is fixed in the simulated dataset without compromising transferability.
The spectral width $\sigma(\omega,T)$ is modeled to reflect realistic energy and temperature dependencies, ensuring alignment with experimentally observed values (in eV and Kelvin):
@@
These dependencies account for electronic self-energy within a Fermi-liquid framework [1] and additional broadening due to phase-mode scattering above $T_{c}$[5, 4, 6]. To enhance transferability across different superconducting materials and experimental conditions, the resulting spectra, after applying the theoretical broadening in Eq. (5), undergo convolution with a resolution function, whose parameters are randomly sampled. The energy-momentum resolution is implemented as a 2D Gaussian convolution on the 2D spectral function, with a typical momentum resolution of 10-pixel size (see Note S4).
@@
The background is treated as 10 times the average per-pixel intensity across the simulated cut multiplied by the Fermi function, which mimics the momentum-scrambling secondary scattering process in photoemission. Since we are mostly in intermediate to high count rate regime, the noise distribution is approximated as a Gaussian with a standard deviation of $\alpha\sqrt{N_{ij}}$, where $N_{ij}$ is the simulated count at pixel $\{i,j\}$, and $\alpha$ is an input parameter to control the signal to noise ratio. This approximation mimics the Poisson noise observed in real experiments, which is generally in the high-count-rate regime and asymptotically approaches a Gaussian noise. The entire spectrum takes an absolute value to eliminate negative counts due to the application of Gaussian noise at extremely low count regimes, which has negligible impact on the main spectral region of interest. A comparison of the parameter space of simulated and experimental data can be found in Note S3. A total of 1,745 simulated spectra were obtained, with 80% (1,395) being utilized as training data.
IV.2 Domain adaptation
@@
DANN is a domain adaptation technique designed to reduce the domain shift between the source (simulated) and target (experimental) domains. The model architecture comprises three parts as shown in Fig. 2: the feature extractor ($G_{f}(\cdot;\theta_{f})$) with parameters $\theta_{f}$, the phase classifier ($G_{y}(\cdot;\theta_{y}$)) with parameters $\theta_{y}$, and the domain classifier ($G_{d}(\cdot;\theta_{d}$)) with parameters $\theta_{d}$. Training DANN involves minimizing the phase classification loss using labels from the source domain while simultaneously encouraging the feature extractor to learn domain-invariant features through adversarial training (see detailed illustrations of forward and back propagation flow in Figure S6). This is achieved using a gradient reversal layer (GRL) (denoted as $\mathcal{R}(\cdot)$), placed between the feature extractor and the domain classifier. The GRL acts as an identity transformation during forward propagation but flips the sign of $\partial\mathcal{L}_{d}/\partial\theta_{f}$ during backpropagation, effectively making the feature extractor maximize the domain classification loss while the domain classifier minimizes it. We define the $\mathcal{L}_{y}$ and $\mathcal{L}_{d}$ as the corresponding cross-entropy loss [38] for label prediction and domain classification, respectively. The general form of the cross-entropy loss for a classification task is:
$\mathcal{L}(p,y)=-\sum_{k=1}^{K}y_{k}\log p_{k}\,,$ (6)
where $K$ is the number of classes; $p$ is the predicted probability vector for each class; $y$ is the one-hot encoded true label vector. Applying this to our specific losses, the phase classification loss for a single sample with true phase label $y\in\{0,1\}$ and predicted probability $\bar{p}_{\textrm{SC}}$ is defined as:
$\mathcal{L}_{y}(\bar{p}_{\textrm{SC}},y)=-y\log(\bar{p}_{\textrm{SC}})-(1-y)\log(1-\bar{p}_{\textrm{SC}})\,.$ (7)
For domain labels $d\in\{0,1\}$ (with $d=0$ for the source domain and $d=1$ for the target domain), and with $\bar{p}_{\textrm{target}}$ being the predicted probability that a sample is from the target domain, the domain classification loss is defined as:
$\mathcal{L}_{d}(\bar{p}_{\textrm{target}},d)=-d\log(\bar{p}_{\textrm{target}})-(1-d)\log(1-\bar{p}_{\textrm{target}})\,.$ (8)
The training objective for the whole DANN is formulated as follows:
[…]
@@
where $\mathbb{E}$ represents the expected value; $\mathbf{x}_{i}^{s}$ denotes a simulated spectrum with phase label $y_{i}^{s}$ from the source domain $D_{s}=\{(\mathbf{x}_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}}$; $\mathbf{x}_{j}^{t}$ denotes an experimental spectrum in the target domain $D_{t}=\{(\mathbf{x}_{j}^{t})\}_{j=1}^{n_{t}}$; $d^{s}$ and $d^{t}$ are the domain labels. Here, $\lambda$ is the adaptation parameter, balancing the two objectives that shape the features during training. Upon training completion, the phase classifier is capable of predicting labels for both source and target domain samples.
IV.3 Model architecture and implementation
@@
Our optimized model employs four convolutional layers with hidden channels of 16, 32, 64, and 64, respectively. Each layer utilizes a 3$\times$3 convolutional kernel, a stride of 1, and padding of 2, followed by max pooling with the same kernel size and a stride of 2, and an activation function. All the activation functions used in this network are rectified linear units (ReLU) except for the first convolutional layer. To alleviate the dying ReLU problem where some neurons become permanently inactive and only output 0 for any input during training, the first convolutional layer ReLU is replaced with the leaky ReLU (LeakyReLU) [39] using a default constant slope of 0.01. Post-convolution, the network applies adaptive average pooling to the feature maps, which are then flattened to interface with the subsequent fully connected layers. The domain classifier and phase classifier are both two-layer fully connected neural networks with 64 hidden dimensions each. To enable phase label prediction, a softmax layer is added to the phase classifier to convert the predicted logits $\mu_{\alpha}$ for each phase $\alpha$ ($\alpha$ = SC or normal) into probability values $p_{\alpha}\propto e^{\mu_{\alpha}}$, obeying the sum rule $\sum_{\alpha}p_{\alpha}=1$. All models are implemented using PyTorch 1.12.1 and trained on a single NVIDIA RTX A5000 GPU. The optimized training procedure employs the Adam optimizer [40] with a learning rate of 0.0005, weight decay (L2 penalty, 0.001), and early stopping. Dropout with a probability of 0.4 is applied after the last convolutional layer and in every layer of each fully connected neural network. Training is performed with mini-batches of 4 spectra for 150 epochs. The adaptation parameter is set to be 1.2 and is gradually increased from 0 through:
@@
IV.4 Model evaluation and hyperparameter tuning
Evaluating the performance of an unsupervised domain adaptation (UDA) model is challenging due to the lack of the target domain labels [41, 42, 43]. As the domain classifier acts as a regularizer on CNN [19], we first optimize the hyperparameter of a CNN model without the domain classifier using simulated ARPES data (detailed in supplemental methods, Figure S3 and Table S5). We then fine-tune the adaptation parameter based on the transfer score (TS) metrics [43], which evaluates the transferability and discriminability of the feature space. The TS is formulated as follows: